OpenStudy (pphalke):

There were 325 tickets sold for a basketball game. The activity cardholders tickets cost $1.50 and the non card holders ticket cost $2.25. The total amount of money collected was $525.00. How many of each kind of ticket were sold?

OpenStudy (giantrobot):

Okay you have to use simultaneous equations. The first equation is simple. Let Cardholder's = x and non Cardholder's = y Therefore \[x +y=325\] The second formula can be derived from the costs. \[1.5x + 2.25y=525\] Now combine the formulas and solve for the x and y one at a time. Here's how I would do it. \[487.5-1.5y+2.25y=525\] \[y=48\] Logic dictates that \[x=277\]